OpenStudy (unimatix):

A 1 kg mass traveling at 2 m/s collides head on with a 3 kg mass traveling at 4 m/s. What is the magnitude of the final momentum (in kg m/s) of the 3 kg mass if the one kg mass rebounds in the opposite direction at 2 m/s?

OpenStudy (unimatix):

Not sure if I'm doing this right. 1kg * 2 m/s = 2 kg m/s 3kg * 4 m/s = 12 kg m/s Total momentum initial = Total momentum final 12 kg m/s = 12 kg m/s So... 1kg * 2 m/s = 2 kg m/s 12kg m/s - 2kg m/s = 10 kg m/s Is 10 kg m/s the right answer? Or is that wrong and I need to add the 2kg m/s in that it is moving the opposite direction. HELP.

OpenStudy (phi):

total momentum of the "system" i.e. the two objects is \[ m_1 \cdot v_1 + m_2\cdot v_2\] based on the question, we know the velocities are opposite directions so we can let one be negative and the other positive

OpenStudy (unimatix):

So I'm thinking now that the initial momentum is -10.

OpenStudy (unimatix):

Which would make the final momentum -10?

OpenStudy (unimatix):

And then if the 1 kg mass has a momentum of -2 Ns. The other will have a momentum of -8 Ns. Which is a magnitude of 8?

OpenStudy (phi):

so you can say 1*-2 + 3*4 = -2 +12=10 (which velocity is negative is arbitrary)

OpenStudy (unimatix):

Makes sense

OpenStudy (unimatix):

That should essentially be the same right?

OpenStudy (phi):

momentum is conserved, so we can say 1*2 + 3*x = 10 we use +2 because we are going the opposite way of -2) 3x=8 magnitude of the final momentum is | 8| = 8

OpenStudy (unimatix):

Sweet awesome. You were very helpful!

OpenStudy (phi):

if we started with 1*2+3 *-4 = -10 1*-2 + 3x= -10 3x= -8 |-8| i.e. the momentum = 8

OpenStudy (unimatix):

Yep. I was thinking they would mathematically end up the same either way. Thanks again!